# RECAST: Reparameterized, Compact weight Adaptation for Sequential Tasks

### v1

@@ A.1 Dataset
CIFAR100  (Krizhevsky, 2009) 100 60000
CIFAR10  (Krizhevsky, 2009) 10 60000
@@ A.3 Reconstruction Analysis
RECAST can reconstruct models across various architecture and scales. In Table 6, we present the results of reconstructed models and the official performance of these models on the Imagenet-1K (Deng et al., 2009) dataset. All reconstructed models in the table have the same configuration - 6 groups, 2 layers in each group and 2 sets of coefficients for each target module (except ViT-Large has 12 groups). The inference comparison on Imagenet tells us that, RECAST can completely emulate the feature generation of any pretrained weight. We also report the averaged best accuracy@top1 across the datasets in Table 3, for the varying scales of the ViT models. In all cases, RECAST perfermos slightly better than baseline without coefficient-tuning, and significantly better with coefficient-tuning.
@@
The Frobenius norm helps ensure diversity among $T_{g,i}$, allowing for a wide range of possible $W_{l,m}$. A higher score means more difference among the template. The entropy ensures that each $T_{g,i}$ is itself balanced and flexible, contributing to the adaptability of the reconstructed weights. A higher entropy corresponds to better information quality within each group. Together, these metrics provide a comprehensive view of the template characteristics in RECAST, helping to optimize both the diversity of the template bank and the flexibility of individual templates for efficient and effective weight reconstruction. In Figure 8(a) and  8(b), we show the change of Frobenius norm and Entropy. The VAE models perform very low in terms of the Frobenius metric, although their pattern varies depending on the loss function. This is also true for the models generated through Neural Mimicry 2.2,i.e. models reconstructed with MSE loss and SmoothL1 loss performance almost similar on classification task but differs in entropy pattern and Frobenius norm. Another observation is that noise perturbations don’t seem to impact these two metrics at all.
@@ A.4.2 RECAST Integration with Existing Methods
RECAST complements existing incremental learning methods as a task-adaptable weight generation backbone. For CNN architectures, RECAST generates base weights that PiggyBack’s (Mallya & Lazebnik, 2018) binary masks can modulate and provides primary convolution weights that CLR’s Ge et al. (2023) Ghost modules can refine. In transformer architectures, RECAST produces core weight matrices that methods like MeLO, AdaptFormer (for query/value and MLP components), DoRA (for magnitude-direction decomposition), and RoSA (for sparse adaptations) can then modify.

### v2

@@ A.1 Dataset
CIFAR100  (Krizhevsky & Hinton, 2009) 100 60000
CIFAR10  (Krizhevsky & Hinton, 2009) 10 60000
@@ A.3 Reconstruction Analysis
RECAST can reconstruct models across various architecture and scales. In Table 6, we present the results of reconstructed models and the official performance of these models on the Imagenet-1K (Deng et al., 2009) dataset. All reconstructed models in the table have the same configuration - 6 groups, 2 layers in each group and 2 sets of coefficients for each target module (except ViT-Large has 12 groups). The inference comparison on Imagenet tells us that, RECAST can completely emulate the feature generation of any pretrained weight. We also report the averaged best accuracy@top1 across the datasets in figure 4.1, for the varying scales of the ViT models. In all cases, RECAST perfermos slightly better than baseline without coefficient-tuning, and significantly better with coefficient-tuning.
@@
The Frobenius norm helps ensure diversity among $T_{g,i}$, allowing for a wide range of possible $W_{l,m}$. A higher score means more difference among the template. The entropy ensures that each $T_{g,i}$ is itself balanced and flexible, contributing to the adaptability of the reconstructed weights. A higher entropy corresponds to better information quality within each group. Together, these metrics provide a comprehensive view of the template characteristics in RECAST, helping to optimize both the diversity of the template bank and the flexibility of individual templates for efficient and effective weight reconstruction. In Figure 9(a) and  9(b), we show the change of Frobenius norm and Entropy. The VAE models perform very low in terms of the Frobenius metric, although their pattern varies depending on the loss function. This is also true for the models generated through Neural Mimicry 3.2,i.e. models reconstructed with MSE loss and SmoothL1 loss performance almost similar on classification task but differs in entropy pattern and Frobenius norm. Another observation is that noise perturbations don’t seem to impact these two metrics at all.
@@ A.4.2 RECAST Integration with Existing Methods
RECAST complements existing incremental learning methods as a task-adaptable weight generation backbone. For CNN architectures, RECAST generates base weights that PiggyBack’s (Mallya et al., 2018) binary masks can modulate and provides primary convolution weights that CLR’s Ge et al. (2023) Ghost modules can refine. In transformer architectures, RECAST produces core weight matrices that methods like MeLO, AdaptFormer (for query/value and MLP components), DoRA (for magnitude-direction decomposition), and RoSA (for sparse adaptations) can then modify.